Use elimination to solve????
x/2+2y=9
2x-y=18

Answer by checkley71(8403)

x/2+2y=9
2x-y=18 multiply by 2
4x-2y=36
x/2+2y=9 adding these two equatons
4 1/2x=45 or x=10
substituting 10 for x in the first equation we get
10/2+2y=9 or 5+2y=9 or 2y=4 or y=2